2x+2y= 2/3
y=3x-1

Answer by Cromlix(4381)

2x + 2y = 2/3...1
y = 3x -1....2
Sort out (2)
2x + 2y = 2/3....1
-3x + y = -1....2
Multiply (1) by 3
Multiply (2) by 2
6x + 6y = 2...3
-6x + 2y = -2...4
Add (3) + (4)
8y = 0
y = 0
Substitute y = 0 in
equation (2)
-3x + y = -1
-3x + 0 = -1
-3x = -1
x = 1/3
x = 1/3 and y = 0
Hope this helps.
;-)